# SiT: Exploring Flow and Diffusion-based Generative Models with Scalable Interpolant Transformers

### v1

@@ 2.6 Interpolant Transformer Architecture
The backbone architecture and capacity of generative models are also crucial for producing high-quality samples. In order to eliminate any confounding factors and focus on our exploration, we strictly follow the standard Diffusion Transformer (DiT) [48] and its configurations. This way, we can also test the scalability of our model across various model sizes.
@@
Similarly, SiT is also a latent generative model and we use the same pre-trained VAE encoder and decoder models originally used in Stable Diffusion [51]. SiT processes a spatial input $z$ (shape $32\times 32\times 4$ for $256\times 256\times 3$ images) by first ‘patchifying’ it into $T$ linearly embedded tokens of dimension $d$. We always use a patch size of 2 in these models as they achieve the best sample quality. We then apply standard ViT [21] sinusoidal positional embeddings to these tokens. We use a series of $N$ SiT transformer blocks, each with hidden dimension $d$.
Our model configurations—SiT-{S,B,L,XL}—vary in model size (parameters) and compute (flops), allowing for a model scaling analysis. For class-conditional generation on ImageNet, we use the AdaLN-Zero block [48] to process additional conditional information (times and class labels). SiT architectural details are listed in Table 3.
Model Layers $N$ Hidden size $d$ Heads
SiT-S 12 384 6
SiT-B 12 768 12
SiT-L 24 1024 16
SiT-XL 28 1152 16
@@ 3 Experiments
To provide a more detailed answer to the question raised in Tab. 1 and make a fair comparison between DiT and SiT, we gradually transition from a DiT model (discretized, score prediction, VP interpolant) to a SiT model (continuous, velocity prediction, Linear interpolant) in the following four subsections, and present the impacts on performance. Throughout our experiments in each subsection, we uses a DiT-B model at 400K training steps as our backbone. For solving the ODE (2), we adopt a fixed Heun integrator; for solving the SDE (4), we used an Euler-Maruyama integrator. With both solver choices we limit the number of function evaluations (NFE) to be $250$ to match the number of sampling steps used in DiT. All numbers presented in the following sections are FID-50K scores evaluated on the ImageNet256 training set.
3.1 Discrete to Continuous Time
To understand the role of continuous-time versus discrete-time models, we study discrete-time DDPM against continuous-time SBDM-VP with estimation of the score. The results are presented in Table 4, where we find a marginal improvement in FID scores when going from a discrete-time denoiser to a continuous-time score field.
@@ 3.1 Discrete to Continuous Time
Table 4: DDPM vs. SBDM.
3.2 Model parameterizations
To clarify the role of the model parameterization in the context of SBDM-VP, we now compare learning (i) a score model using (6), (ii) a weighted score model (see Sec. A.3), or (iii) a velocity model using (7). The results are shown in Table 5, where we find that we obtain a significant performance improvement by learning a weighted score model or a velocity model.
Interpolant Model Objective ODE
@@ 3.2 Model parameterizations
Table 5: Effect of model parameterization.
3.3 Choice of interpolant
Figure 4: Path length. The path length $\mathcal{C}(v)=\int_{0}^{1}\mathbb{E}[|{\mathbf{v}}({\mathbf{x}}_{t},t)|^{2}]{\mathrm{d}}t$ arising from the velocity field at different training steps for the various models considered: SBDM (VP), Linear, and GVP; each curve is approximated by $10000$ datapoints.
Section 2 highlights that there are many possible ways to build a connection between the data distribution and a Gaussian by varying the choice of $\alpha_{t}$ and $\sigma_{t}$ in the definition of the interpolant (1). To understand the role of this choice, we now study the benefits of moving away from the commonly-used SBDM-VP setup. We consider learning a velocity model ${\mathbf{v}}({\mathbf{x}},t)$ with the Linear and GVP interpolants presented in (12), which make the interpolation between the Gaussian and the data distribution exact on $[0,1]$. We benchmark these models against the SBDM-VP in Table 6, where we find that both the GVP and Linear interpolants obtain significantly improved performance.
@@ 3.3 Choice of interpolant
Table 6: Effect of interpolant.
3.4 Deterministic vs stochastic sampling
Using the SBDM diffusion coefficient.
As shown in Sec. 2, given a learned model, we can sample using either the probability flow equation (2) or an SDE (4). In SBDM we conventionally take as diffusion coefficient $w_{t}=\beta_{t}$. For Linear and GVP interpolant, we follow the derivation in Appendix B to express $\beta_{t}$ in terms of $\alpha_{t}$ and $\sigma_{t}$.
Our results are shown in Tab. 7, where we find performance improvements by sampling with an SDE over the ODE, which is in line with the bounds given in [3]: the SDE has better control over the KL divergence between the time-dependent density at $t=0$ and the ground truth data distribution. We note that the performance of ODE and SDE integrators may differ under different computation budgets. As shown in Fig. 5, the ODE converges faster with fewer NFE, while the SDE is capable of reaching a much lower final FID score when given a larger computational budget.
@@ 3.4 Deterministic vs stochastic sampling
Table 7: ODE vs. SDE, SBDM diffusion.
@@
Motivated by the improved performance of SDE sampling, we now consider the effect of tuning the diffusion coefficient in a manner that is distinct from the choices made in SBDM, as detailed in Tab. 2. As shown in Tab. 8, we find that the optimal choice for sampling is both model prediction and interpolant dependent. We picked the three functions with the best performance and sweep through all different combinations of our model prediction and interpolant, and present the result in Tab. 8.
We also note that the influences of different diffusion coefficients can vary across different model sizes. Empirically, we observe the best choice for our SiT-XL is a velocity model with Linear interpolant and sampled with $\sigma_{t}$ coefficient.
Interpolant Model Objective $w_{t}=\beta_{t}$ $w_{t}=\sigma_{t}$ $w_{t}=\sin^{2}(\pi t)$
SBDM-VP velocity $\mathcal{L}_{\mathrm{v}}$ 37.8 38.7 39.2
score $\mathcal{L}_{\mathrm{s}_{\lambda}}$ 35.7 37.1 37.7
GVP velocity $\mathcal{L}_{\mathrm{v}}$ 32.9 33.4 33.6
score $\mathcal{L}_{\mathrm{s}}$ 38.0 33.5 33.2
Linear velocity $\mathcal{L}_{\mathrm{v}}$ 33.6 33.5 33.3
score $\mathcal{L}_{\mathrm{s}}$ 41.0 35.3 34.4
Table 8: Evaluation of our SDE samplers. All results in the table are FID-50K scores produced by an SiT-B model at 400K training steps. The last three columns specify different diffusion coefficients $w_{t}$ detailed in Tab. 2. To make the SBDM-VP competitive when learning a score, we use a weighted score given in Sec. A.4, as per the remarks below (7).
@@ 3.5 Classifier-free guidance
Guidance for a velocity field means that: (i) that the velocity model ${\mathbf{v}}_{\theta}({\mathbf{x}},t;{\mathbf{y}})$ takes class labels $y$ during training, where $y$ is occasionally masked with a null token $\emptyset$; and (ii) during sampling the velocity used is ${\mathbf{v}}_{\theta}^{\zeta}({\mathbf{x}},t;{\mathbf{y}})=\zeta{\mathbf{v}}_{\theta}({\mathbf{x}},t;{\mathbf{y}})+(1-\zeta){\mathbf{v}}_{\theta}({\mathbf{x}},t;\emptyset)$ for a fixed $\zeta>0$. In Appendix C, we show that this indeed corresponds to sampling the tempered density $p({\mathbf{x}}_{t})p({\mathbf{y}}|{\mathbf{x}}_{t})^{\zeta}$ as proposed in [46]. Given this observation, one can leverage the usual argument for classifier-free guidance of score-based models.
For a CFG scale of $\zeta=1.5$, DiT-XL sees an improvement in FID from 9.6 (non-CFG) down to 2.27 (CFG). We observed similar performance improvement with our largest SiT-XL model under identical computation budget and CFG scale. Sampled with an ODE, the FID-50K score improves from 9.4 to 2.15 (Tab. 9 and Appendix D); with an SDE, the FID improves from 8.6 to 2.06 (Tab. 1 and Tab. 9). This shows that SiT benefits from the same training and sampling choices explored previously, and can surpass DiTs performance in each training setting, not only with respect to model size, but also with respect to sampling choices.
Class-Conditional ImageNet 256x256
@@
SiT-XL(cfg = 1.5, SDE:$\sigma_{t}$) 2.06 4.50 270.27 0.82 0.59
Table 9: Benchmarking class-conditional image generation on ImageNet 256x256. SiT-XL surpasses DiT-XL in FID when either of the samplers, ODE or SDE-based.
@@ 4 Related Work
The transformer architecture [66] has emerged as a powerful tool for application domains as diverse as vision [47, 21], language [69, 68], quantum chemistry [67], active matter systems [9], and biology [11]. Several works have built on DiT and have made improvements by modifying the architecture to internally include masked prediction layers [22, 70]; these choices are orthogonal to the transition from DiT to SiT studied in this work; they may be fruitfully combined in future work.
@@
Diffusion models arose from [59, 25, 62] and have close historical relationship with denoising methods [57, 29, 28]. Various efforts have gone into improving the sampling algorithms behind these methods in the context of DDPM [60] and SBDM [61, 32]; these are also orthogonal to our studies and may be combined to push for better performance in future work. Improved Diffusion ODE [71] also studies several combinations of model parameterizations (velocity versus noise) and paths (VP versus Linear) for sampling an ODE; they report best results for velocity model with smoother probability flow; they focus on lower dimensional experiments, benchmark with likelihoods, and do not consider SDE sampling. In our work, we explore the effects of changing between VP, Linear, and GVP interpolants, as well as score and velocity parameterizations in depth and show how these choices individually improve performance on the larger scale ImageNet256. We also document how FIDs change with respect to a family of sampling algorithms including black-box ODEs and SDEs indexed by a choice of diffusion coefficients, and show that the best coefficient choice may depend on the model and interpolant. This brings the observations about the flexibility and trade-offs of sampling from [3] into practice.
@@
Velocity field parameterizations using the Linear interpolant were also studied in [39, 42], and were generalized to the manifold setting in [6]. A trade-off in bounds on the KL divergence between the target distribution and the model arises when considering sampling with SDEs versus ODE; [3] shows that minimizing the objectives presented in this work controls KL for SDEs, but not for ODEs. Error bounds for SDE-based sampling with score-based diffusion models are studied in [14, 37, 38, 13]. Error bounds on the ODE are also explored in [15, 7], in addition to the Wasserstein bounds provided in [1].
Other related works make improvements by changing how noise and data are sampled during training. [64, 50] compute mini-batch optimal couplings between the Gaussian and data distribution to reduce the transport cost and gradient variance; [4] instead build the coupling by flowing directly from the conditioning variable to the data for image-conditional tasks such as super-resolution and in-painting. Finally, various work considers learning a stochastic bridge connecting two arbitrary distributions [49, 56, 41, 18]. These directions are compatible with our investigations; they specify the learning problem for which one can then vary the choices of model parameterizations, interpolant schedules, and sampling algorithms.
@@ 5 Conclusion
Acknowledgements.
We would like to thank Adithya Iyer, Sai Charitha Akula, Fred Lu, Jitao Gu, and Edwin P. Gerber for helpful discussions and feedback. The research is partly supported by the Google TRC program.
Appendix
Appendix A Proofs
In all proofs below, we use $\cdot$ for dot product and assume all bold notations (${\mathbf{x}}$, ${\boldsymbol{\varepsilon}}$, etc.) are real-valued vectors in $\mathbb{R}^{d}$. Most proofs are derived from Albergo et al. [3].
A.1 Proof of the probability flow ODE (2) with the velocity in Eq. 3.

### v2

@@ 2.5 Interpolant Transformer Architecture
The backbone architecture and capacity of generative models are both crucial for producing high-quality samples. In order to eliminate any confounding factors and focus on our exploration, we strictly follow the standard Diffusion Transformer (DiT) [50] and its configurations. This way, we can also test the scalability of our model across various model sizes.
@@
Similarly, SiT is a latent generative model, and we use the same pre-trained VAE encoder and decoder models originally used in Stable Diffusion [53]. SiT processes a spatial input $z$ (shape $32\times 32\times 4$ for $256\times 256\times 3$ images) by first ‘patchifying’ it into $T$ linearly embedded tokens of dimension $d$. We always use a patch size of 2 in these models as they achieve the best sample quality. We then apply standard ViT [21] sinusoidal positional embeddings to these tokens. We use a series of $N$ SiT transformer blocks, each with hidden dimension $d$.
Our model configurations—SiT-{S,B,L,XL}—vary in model size (parameters) and compute (flops), allowing for a model scaling analysis. For class-conditional generation on ImageNet, we use the AdaLN-Zero block [50] to process additional conditional information (times and class labels). SiT architectural details are listed in Appendix 0.E.
The complete SiT design space that we explore consists of the choice of time discretization and the model prediction (Sec. 2.2), the choice of the interpolant (Sec. 2.3), the choice of sampler and diffusion coefficient (Sec. 2.4), and the model size (Sec. 2.5).
@@ 3 Experiments
To provide a more detailed answer to the question raised in Table 1 and make a fair comparison between DiT and SiT, we gradually transition from a DiT model (discrete, score prediction, VP interpolant) to a SiT model (continuous, velocity prediction, Linear interpolant) and present the impacts on performance.
Experimental setup.
In the transition experiments, we use SiT-B models trained on $256\times 256$ image resolution on the ImageNet as our backbone. We fix training steps to be 400K throughout the transition. For solving the ODE (2), we adopt a fixed step second-order Heun integrator; for solving the SDE (4), we used a first-order Euler-Maruyama integrator. With both solver choices we limit the number of function evaluations (NFE) to be $250$ to match the number of sampling steps used in DiT. All metrics presented are FID-50K scores evaluated on the ImageNet training set unless otherwise stated.
We also scale up our SiT model to the XL configuration and train on both $256\times 256$ and $512\times 512$ resolution on ImageNet. We strictly follow the training settings of DiT and did not tune any hyperparameters.
3.1 Model Parameterization
Discrete- to continuous-time.
Continuous time training has been previously studied from the perspective of improved likelihood bounds [64, 37]. As mentioned in Section 2.2, here we focus on the fact that training in continuous time allows us to decouple discretization choices in sampling from the particular training method, which allows for finding the right discretization for various choices of diffusion coefficients that we are free to choose after training. We observe a marginal performance increase in Section 3.1 by switching to continuous time.
We additionally observe in Figure 5 that flexibility in integration allows one to trade-off number of functional evaluations and FID performance.
Model parameterization.
To clarify the role of the model parameterization in the context of SBDM-VP, we now compare learning (i) a score model using (6) ($\mathcal{L}_{\mathrm{s}}$), (ii) a weighted score model ($\mathcal{L}_{\mathrm{s}_{\lambda}}$), or (iii) a velocity model using (7)($\mathcal{L}_{\mathrm{v}}$). We observe a significant performance increase with $\mathcal{L}_{\mathrm{s}_{\lambda}}$ and $\mathcal{L}_{\mathrm{v}}$ in Table 3.1.
In accordance with the observation made in [36], we carefully choose a $\lambda(t)$ such that $\lambda_{\mathrm{s}_{\lambda}}$ is made equivalent to $\lambda_{\mathrm{v}}$. We will provide detailed derivations in Appendix 0.A.3, and demonstrate such $\lambda$ is closely related to the maximum likelihood weighting proposed in [63, 66]. Furthermore, we note that $\lambda(t)\to\infty$ as $t\to 0$, thus compensating for the vanishing gradient of the score objective when near the data. This could also account for the performance gain from $\lambda_{\mathrm{s}}$ to $\lambda_{\mathrm{s}_{\lambda}}$.
Table 2: Discrete vs. continuous.
@@ 3.1 Model Parameterization
Table 3: Effect of parameterizations.
Interpolant Model Objective FID
@@
Choices of interpolant.
Sec. 2 highlights that there are many possible ways to build a connection between the data distribution and a Gaussian by varying the choice of $\alpha_{t}$ and $\sigma_{t}$ in the definition of the interpolant (1). To understand the role of this choice, we now study the benefits of moving away from the commonly-used SBDM-VP setup. We consider learning a velocity model ${\mathbf{v}}({\mathbf{x}},t)$ with the Linear and GVP interpolants presented in (12), which make the interpolation between the Gaussian and the data distribution exact on $[0,1]$. We benchmark these models against the SBDM-VP in Table 3.1, where we find that both the GVP and Linear interpolants obtain significantly improved performance.
@@
Table 5: ODE vs. SDE, $w_{t}=w_{t}^{\mathrm{KL}}$.
@@
3.2 Deterministic vs stochastic sampling
As shown in Sec. 2, given a learned model, we can sample using either the probability flow equation (2) or an SDE (4). In Sec. 3.1 we illustrate the discrepancy between the two methods when using the same trained velocity model. We find performance improvements by sampling with an SDE over the ODE, which is in line with the bounds given in [2]: the SDE has better control over the KL divergence between the model density at $t=0$ and the ground truth data distribution. We also note that the performance of ODE and SDE integrators may differ under different computation budgets. As shown in Fig. 5, the ODE converges faster with fewer NFE, while the SDE is capable of reaching a much lower final FID score when given a larger computational budget.
Figure 4: Path length. The path length $\mathcal{C}(v)=\int_{0}^{1}\mathbb{E}[|{\mathbf{v}}({\mathbf{x}}_{t},t)|^{2}]{\mathrm{d}}t$ arising from the velocity field at different training steps; each curve is approximated by $10000$ datapoints at each training step.
@@ 3.2 Deterministic vs stochastic sampling
Motivated by the improved performance of SDE sampling, we now consider the effect of tuning the diffusion coefficient in inference. As shown in Table 6, we sweep through all different combinations of our model prediction and interpolant, and present the result. We find that the optimal choice for sampling is both model prediction and interpolant dependent.
According to Sec. 2.4, the choice of $w_{t}=w_{t}^{\mathrm{KL}}$ would ideally minimize the upper bound for the KL divergence $D_{\mathrm{KL}}(p({\mathbf{x}})|\|p_{0}({\mathbf{x}}))$ and make the SDE approximate the data distribution more closely, barring integration costs. This theoretical result is supported by empirical observation for the SBDM-VP and GVP interpolants presented in Table 6. For Linear interpolants, the cost-regularized version $w_{t}^{\mathrm{KL},\eta}$ provides the best FID, because the SDE for the Linear interpolant with $w_{t}^{\mathrm{KL}}$ becomes hard to integrate at the endpoint. Generally speaking, the score models perform worse than the velocity models, which may be due to the singularity of the objective in (6). Moreover, the efficacy of using $w_{t}^{\mathrm{KL}}$ in this context is also reduced, for similar reason. For example, reverting (9) to obtain $\mathrm{v}_{\theta}({\mathbf{x}},t)$ from $\mathrm{s}_{\theta}({\mathbf{x}},t)$ will result in a singularity at $t=1$ in $\mathcal{L}_{t}$ used to choose $w_{t}^{\mathrm{KL},\eta}$ in (14). Lastly, for SBDM-VP we observe worse result from $w_{t}^{\mathrm{KL},\eta}$ as opposed to $w_{t}^{\mathrm{KL}}$. Different from Linear and GVP, as stated in Sec. 2.4 and Sec. 3.1, $w_{t}^{\mathrm{KL}}$ is well-defined everywhere on $[0,1]$ for SBDM-VP, whereas $w_{t}^{\mathrm{KL},\eta}$ suffers from the singularity issue posed by $\mathcal{L}_{\mathrm{v}}$ near $t=0$. These observations supports our claim made before, that the optimal choice of $w_{t}$ will always be model prediction and interpolant dependent.
We also note that the influences of different diffusion coefficients can vary across different model sizes. Empirically, we observe the best choice for our SiT-XL is a velocity model with Linear interpolant and sampled with $w_{t}^{\mathrm{KL},\eta}$.
Table 6: Evaluation of our SDE samplers. The last three columns specify different diffusion coefficients $w_{t}$. To make the SBDM-VP competitive, we perform evaluation on the weighted score model $\mathcal{L}_{\mathrm{s}_{\lambda}}$. We mark the optimal $w_{t}$ for each interpolant.
Interpolant Model Objective $w_{t}=w_{t}^{\mathrm{KL}}$ $w_{t}=\sigma_{t}$ $w_{t}=\sin^{2}(\pi t)$ $w_{t}=w_{t}^{\mathrm{KL},\eta}$
SBDM-VP velocity $\mathcal{L}_{\mathrm{v}}$ 37.8 38.7 39.2 41.1
score $\mathcal{L}_{\mathrm{s}_{\lambda}}$ 35.7 37.1 37.7 38.9
GVP velocity $\mathcal{L}_{\mathrm{v}}$ 32.9 33.4 33.6 33.2
score $\mathcal{L}_{\mathrm{s}}$ 37.8 33.5 33.2 33.3
Linear velocity $\mathcal{L}_{\mathrm{v}}$ 33.6 33.5 33.3 33.0
score $\mathcal{L}_{\mathrm{s}}$ 41.0 35.3 34.4 34.9
@@ 3.3 Classifier-free guidance
Guidance for a velocity field means that: (i) that the velocity model ${\mathbf{v}}_{\theta}({\mathbf{x}},t;{\mathbf{y}})$ takes class labels $y$ during training, where $y$ is occasionally masked with a null token $\emptyset$; and (ii) during sampling the velocity used is ${\mathbf{v}}_{\theta}^{\zeta}({\mathbf{x}},t;{\mathbf{y}})=\zeta{\mathbf{v}}_{\theta}({\mathbf{x}},t;{\mathbf{y}})+(1-\zeta){\mathbf{v}}_{\theta}({\mathbf{x}},t;\emptyset)$ for a fixed $\zeta>0$. In Appendix 0.C, we show that this indeed corresponds to sampling the tempered density $p({\mathbf{x}}_{t})p({\mathbf{y}}|{\mathbf{x}}_{t})^{\zeta}$ as proposed in [48]. Given this observation, one can leverage the usual argument for classifier-free guidance of score-based models.
We observed similar performance improvement with our SiT-XL models under identical computation budget and CFG scale as DiT-X: models. For SiT-XL $256\times 256$, we follow identical settings in DiT and train the model for 7M steps. We show samples in Fig. 1, and report the result in Table 7. For SiT-XL $512\times 512$, we train the model for 3M steps under the same setting and report the result in Table 7. Under both training settings we observe performance advantage of SiT. We display more samples in Fig. 1 and in Appendix 0.F.
Table 7: Benchmarking class-conditional image generation on ImageNet $256\times 256$ and $512\times 512$. SiT-XL surpasses DiT-XL in both resolutions.
Class-Conditional ImageNet $256\times 256$
@@
SiT-XL(cfg = 1.5, SDE) 2.06 4.49 277.50 0.83 0.59
Class-Conditional ImageNet $512\times 512$
[…]
BigGAN-deep[10] 8.43 8.13 177.90 0.88 0.29
StyleGAN-XL[57] 2.41 4.06 267.75 0.77 0.52
Mask-GIT[12] 7.32 - 156.0 - -
ADM[19] 23.24 10.19 58.06 0.73 0.60
ADM-G, ADM-U 3.85 5.86 221.72 0.84 0.53
Simple Diffusion(U-Net)[28] 4.28 - 171.0 - -
Simple Diffusion(U-ViT, L) 4.53 - 205.3 - -
VDM++[36] 2.65 - 278.1 - -
DiT-XL(cfg = 1.5)[50] 3.04 5.02 240.82 0.84 0.54
SiT-XL(cfg = 1.5, SDE) 2.62 4.18 252.21 0.84 0.57
@@ 4 Related Work
The transformer architecture [67] has emerged as a powerful tool for application domains as diverse as vision [49, 21], language [69, 68], quantum chemistry [23], active matter systems [9], and biology [11]. Several works have built on DiT and have made improvements by modifying the architecture to internally include masked prediction layers [22, 70]; these choices are orthogonal to this work and may be fruitfully combined in future work.
@@
Diffusion models arose from [61, 25, 64] and have close historical relationship with denoising methods [59, 30, 29]. Various efforts have gone into improving the sampling algorithms behind these methods in the context of DDPM [62] and SBDM [63, 33]; these are also orthogonal to our studies and may be combined to push for better performance in future work. Improved Diffusion ODE [71] also studies several combinations of model parameterizations (velocity versus noise) and paths (VP versus Linear). Unlike our work, they focus on lower dimensional experiments, benchmark with likelihoods, and do not consider SDE sampling.
@@
Velocity parameterizations using the Linear interpolant were also studied in [41, 43], and were generalized to the manifold setting in [6]. A trade-off in bounds on the KL divergence between the target distribution and the model arises when considering sampling with SDEs versus ODE; [2] shows that minimizing the objectives presented in this work controls KL for SDEs, but not for ODEs. Error bounds for SDE-based sampling with score-based diffusion models are studied in [14, 38, 39, 13], for ODE-base sampling are explored in [15, 7], in addition to the Wasserstein bounds provided in [4].
Other related works make improvements by changing how noise and data are sampled during training. [65, 52] compute mini-batch optimal couplings between the Gaussian and data distribution to reduce the transport cost and gradient variance; [3] instead build the coupling by flowing directly from the conditioning variable to the data for image-conditional tasks. Finally, various work considers learning a stochastic bridge connecting two arbitrary distributions [51, 58, 44, 18]. These directions are compatible with our investigations; they specify the learning problem for which one can vary the choices of model parameterizations, interpolant schedules, and sampling algorithms.
@@ 5 Conclusion
Acknowledgements
We would like to thank Adithya Iyer, Sai Charitha Akula, Fred Lu, Jiatao Gu, and Edwin P. Gerber for helpful discussions and feedback. The research is partly supported by the Google TRC program.
Appendix 0.A Proofs
In all proofs below, we use $\cdot$ for dot product and assume all bold notations (${\mathbf{x}}$, ${\boldsymbol{\varepsilon}}$, etc.) are real-valued vectors in $\mathbb{R}^{d}$. Most proofs are derived from [2].
0.A.1 Proof of the probability flow ODE (2) with the velocity in Eq. 3.